# Model Tampering Attacks Enable More Rigorous Evaluations of LLM Capabilities

### v1

@@ 4.2 Jailbreaking
We repeat analogous experiments with similar results in the jailbreaking setting. Finally, to test the generality of our findings outside the unlearning paradigm, we ask whether they extend to jailbreaking. Using the 9 off-the-shelf models and 11 attacks from Table 1, we conducted parallel experiments as in Section 4.1 but with attacks that jailbreak models for compliance with harmful requests. We plot all results in Appendix D.
@@
We then reproduced our analysis of whether the success of model tampering jailbreaks correlates with and/or upper-bound the success of input-space jailbreaks (Figure 17). Like before, we find that fine-tuning attack success tends to empirically upper bound the success of input-space attacks. However, unlike before, we do not find clear evidence of a reliable correlation between tampering and input-space attacks due to only having 9 samples.
@@ 5 Discussion
Implications for evaluations: Our findings have direct implications for AI risk evaluations and governance. Current evaluation frameworks rely heavily on input-space attacks, but our results show that these methods systematically underestimate worst-case behavior. Model tampering attacks provide a powerful tool for identifying hidden risks. By modifying a model’s internals—either through activation perturbations or fine-tuning—we can reveal failure modes that input-space attacks miss. This is particularly critical for open-weight models, where safety mitigations can be undone post-release. Moreover, our findings on the fragility of unlearning reversibility highlight another governance concern. If harmful capabilities can be reactivated with minimal tampering, then regulators and developers must reconsider claims that unlearning provides robust safety assurances.
@@
  • Models and tasks: Experiments with larger models and more tasks would offer a broader understanding of model tampering attacks’ potential to aid in evaluations.
@@ Acknowledgments
We are grateful to the Center for AI Safety for compute and the Machine Learning Alignment and Theory Scholars program for research support. We thank Daniel Filan for technical support, and Antony Kellermann for technical discussion. Finally, were are grateful to Peter Henderson, Ududec Cozmin, Ekdeep Singh Lubana, and Taylor Kulp-McDowall for their feedback.
@@ A.4 Attack Methods and Implementation
To test each model’s performance on benign pruning attacks, we use Weights and Activations (WandA) pruning (Sun et al., 2023). WandA prunes the lowest $p\%$ of parameters according to the importance score generated by $W_{i}X$, where $W_{i}$ is the $i$th layer weight and $X$ is the input. For all experiments, we prune the weights 15%.
@@ B.1 Standard Attacks
In Figure 8, we plot the unlearning attack successes for all model tampering attacks against all input-space attacks.
@@ B.2 UK AISI Attacks and Evaluation
In Figure 9, we plot the full unlearning attack successes for all model tampering attacks against the UK AISI attack.
Figure 9: Model tampering attacks remain predictive for a proprietary attack from the UK AI Safety Institute. (a) In these experiments, correlations are weaker than with non- UK AISI attacks, but benign fine-tuning attacks continue to correlate with UK AISI input-space attack success. (b) Fine-tuning attacks still tend to exceed the success of input-space attacks, though less consistently than with the attacks from Figure 5.
Next, to test the limits of our hypothesis that model tampering attacks can help evaluators assess unforeseen failure modes, we evaluated model performance under an entirely different non-WMDP benchmark for dual-use bio capabilities from the UK AI Safety Institute. Figure 11 shows that WMDP-Bio performance correlates with this evaluation with $r=0.64$ and $p=0.0$. To correct for this confounding factor, in Figure 6, we use model tampering attack success on WMDP-Bio to predict the residuals from a linear regression predicting UK AISI Bio evaluation results from WMDP-Bio evaluation results. Here, we find weak correlations except for the case of the pruning and benign fine-tuning methods. Overall, this suggests that while model tampering attacks can be informative about unforeseen failure modes across different attacks, they do not necessarily do so across different tasks.
@@ C.1 Model tampering attacks improve predictive accuracy for worst-case input-space vulnerabilities
Figure 13: Model tampering attacks help predict worst-case input-space vulnerabilities. We perform linear regressions to predict the worst-case input-space success rate from success rates of subsets of attacks. Including model tampering attacks in these subsets improves worst-case vulnerability estimation $R^{2}$ by 0.05-0.1.
In this section, we investigate the utility of model tampering attacks for worst-case input-space vulnerability estimation. While Figure 5 shows that fine-tuning attacks empirically upper bound worst-case input-space vulnerabilities, in this section, we also show that model tampering attacks improve evaluators’ ability to predict worst-case vulnerabilities – even if they already have access to input-space attacks.
For all experiments in this section, we assume the setting of an evaluator who only has access to a subset of attacks in order to estimate worst-case input-space vulnerabilities (potentially due to unforeseen attacks). Whether due to resource constraints on the number of evaluations that are feasible to implement or due to the constant invention of new attack methods, evaluators will always be in this kind of setting. In our setup, we fit linear regression to predict worst-case input-space success rates given the success rates of a subset of attacks. Our dataset consists of a table of all unlearned models (and their 8 checkpoints throughout training) and all attack success scores (WMDP accuracy after attack - base WMDP accuracy). We perform $k$-fold cross-validation across model families by holding out all models trained by the same unlearning method, one method at a time. We then average statistics (e.g. RMSE, $R^{2}$) across the splits. Note that we include the UK AISI input-space attack in these experiments, giving us 6 input-space attacks.
@@ Appendix D Full Jailbreaking Results
Figure 17: In our experiments, fine-tuning attack successes empirically exceed the successes of state-of-the-art input-space attacks for jailbreaking. Here, we plot the increases in compliance with harmful requests under model tampering attacks against the best-performing (out of 5) input-space attacks for each model. On the right, the $x$ axis is the best (over 2) between a LoRA and Full fine-tuning attack. We also display the correlation and the correlation’s $p$ value. There are only 9 points in each figure, so we cannot draw strong conclusions. However, we see no clear evidence of a correlation between model tampering and input-space attack success. However, as before in Figure 5, fine-tuning attacks empirically tend to upper bound the success of input-space attacks. The only case out of 9 in which this was not the case was with the uncensored Orenguteng model (link) which was unlike the other 8 in that it was not designed to be robust to jailbreaks.
@@
Figure 19: Single-turn model tampering attack successes correlate with attacks from Cascade, a multi-turn, proprietary attack algorithm . Since Cascade is state-of-the-art and multi-turn, our single-turn model tampering attacks do not tend to empirically upper-bound the success of this attack as we find for unlearning experiments (Figure 5). However, they empirically correlate with its success.

### v2

@@ 4.2 Jailbreaking Experiments
We repeat analogous experiments with similar results in the jailbreaking setting. Finally, to test the generality of our findings outside the unlearning paradigm, we ask whether they extend to jailbreaking. Using the 9 off-the-shelf models and 11 attacks from Table 1, we conducted parallel experiments as in Section 4.1 but by pitting off-the-shelf refusal-finetuned models against jailbreaking attacks. We plot all results in Appendix D.
@@
We then reproduced our empirical analysis of whether the success of model tampering jailbreaks correlates with and/or conservatively exceeds the success of input-space jailbreaks (Figure 17). Like before, we find that fine-tuning attack success tends to empirically exceed the success of input-space attacks, thus offering a conservative estimation method. However, unlike before, we do not find clear evidence of a reliable correlation between tampering and input-space attacks due to only having 9 samples.
@@ 5 Discussion
Implications for evaluations and safety cases: Our findings have direct implications for performing AI risk evaluations and constructing safety cases (Clymer et al., 2024). Current evaluation frameworks rely heavily on input-space attacks which can easily fail to underestimate worst-case failures. Model tampering attacks provide a useful tool for studying novel, potentially unforeseen risks. By modifying a model’s internal mechanisms — either through activation perturbations or fine-tuning — we can infer the potential existence of failure modes that input-space evaluations may miss (see also Hofstätter et al. (2025)). This is particularly critical for open-weight models, where safety mitigations can be undone post-release.
@@
  • Can models be robust to tampering attacks? This paper and concurrent work (Qi et al., 2024b) show that even defenses designed to make models robust to tampering can be easily undone. We are currently working to better understand tampering robustness and improve the extent to which models can be made robust to tampering attacks.
@@ Acknowledgments
We are grateful to the Center for AI Safety for compute and the Machine Learning Alignment and Theory Scholars program for research support. We thank Daniel Filan for technical support, and Antony Kellermann for technical discussion. Finally, we are grateful to Peter Henderson, Ududec Cozmin, Ekdeep Singh Lubana, and Taylor Kulp-McDowall for their feedback.
@@ A.4 Attack Methods and Implementation
To test each model’s performance on benign pruning attacks, we use Weights and Activations (WandA) pruning Sun et al. (2023). WandA prunes the lowest $p\%$ of parameters according to the importance score generated by $W_{i}X$, where $W_{i}$ is the $i$th layer weight and $X$ is the input. For all experiments, we prune the weights 15%.
@@ B.1 Standard Attacks
In Figure 8, we plot the attack successes for all model tampering attacks against all input-space attacks.
@@ B.2 UK AISI Attacks and Evaluation
In Figure 9, we plot the full attack successes for all model tampering attacks against the UK AISI attack.
Figure 9: Model tampering attacks remain predictive for a proprietary attack from the UK AI Security Institute. (a) In these experiments, correlations are weaker than with non- UK AISI attacks, but benign fine-tuning attacks continue to correlate with UK AISI input-space attack success. (b) Fine-tuning attacks still tend to exceed the success of input-space attacks, though less consistently than with the attacks from Figure 5.
Next, to test the limits of our hypothesis that model tampering attacks can help evaluators assess novel, unforeseen failure modes, we evaluated model performance under an entirely different non-WMDP benchmark for dual-use bio capabilities from the UK AI Security Institute. Figure 11 shows that WMDP-Bio performance correlates with this evaluation with $r=0.64$ and $p=0.0$. To correct for this confounding factor, in Figure 6, we use model tampering attack success on WMDP-Bio to predict the residuals from a linear regression predicting UK AISI Bio evaluation results from WMDP-Bio evaluation results. Here, we find weak correlations except for the case of the pruning and benign fine-tuning methods. Overall, this suggests that while model tampering attacks can be informative about novel failure modes across different attacks, they do not necessarily do so across different tasks.
@@ C.1 Model tampering attacks improve predictive accuracy for worst-case input-space vulnerabilities
Figure 13: Model tampering attacks help predict worst-case input-space vulnerabilities. We perform linear regressions to predict the worst-case input-space success rate from success rates of subsets of attacks. Including model tampering attacks in these subsets improves worst-case vulnerability estimation $R^{2}$ by 0.05-0.1. Ultimately, however, this is likely a conservative quantification of the marginal predictiveness of model tampering attacks for unforeseen input-space threats. The two most effective input space attacks were GCG and AutoPrompt, and as shown in Figure 12, their correlation is 0.88. However, unforeseen attacks in the real world are by no means guaranteed to be as similar to standard input-space attacks as GCG and AutoPrompt are to each other. As a result, this experiment is likely to paint a more pessimistic view on the value of model tampering attacks for predicting held-out input space attacks.
In this section, we investigate the utility of model tampering attacks for worst-case input-space vulnerability estimation. While Figure 5 shows that fine-tuning attacks empirically offer conservative estimates for worst-case input-space vulnerabilities, in this section, we also show that model tampering attacks improve evaluators’ ability to predict worst-case vulnerabilities – even if they already have access to input-space attacks.
For all experiments in this section, we assume the setting of an evaluator who only has access to a subset of attacks in order to estimate worst-case input-space vulnerabilities (potentially due to novel attacks). Whether due to resource constraints on the number of evaluations that are feasible to implement or due to the constant invention of new attack methods, evaluators will always be in this kind of setting. In our setup, we fit linear regression to predict worst-case input-space success rates given the success rates of a subset of attacks. Our dataset consists of a table of all unlearned models (and their 8 checkpoints throughout training) and all attack success scores (WMDP accuracy after attack - base WMDP accuracy). We perform $k$-fold cross-validation across model families by holding out all models trained by the same unlearning method, one method at a time. We then average statistics (e.g. RMSE, $R^{2}$) across the splits. Note that we include the UK AISI input-space attack in these experiments, giving us 6 input-space attacks.
@@ Appendix D Full Jailbreaking Results
Figure 17: In our experiments, fine-tuning attack successes empirically exceed the successes of state-of-the-art input-space attacks for jailbreaking. Here, we plot the increases in compliance with harmful requests under model tampering attacks against the best-performing (out of 5) input-space attacks for each model. On the right, the $x$ axis is the best (over 2) between a LoRA and Full fine-tuning attack. We also display the correlation and the correlation’s $p$ value. There are only 9 points in each figure, so we cannot draw strong conclusions. However, we see no clear evidence of a correlation between model tampering and input-space attack success. However, as before in Figure 5, fine-tuning attacks empirically tend to offer conservative estimates of the success of input-space attacks. The only case out of 9 in which this was not the case was with the uncensored Orenguteng model (link) which was unlike the other 8 in that it was not designed to be robust to jailbreaks.
@@
Figure 19: Single-turn model tampering attack successes correlate with attacks from Cascade, a multi-turn, proprietary attack algorithm . Since Cascade is state-of-the-art and multi-turn, our single-turn model tampering attacks do not tend to empirically exceed the success of this attack as we find for unlearning experiments (Figure 5). However, they empirically correlate with its success.